OpenStudy (anonymous):

What are the coordinates of the center of the ellipse shown below? (x-1)^2/9+(y+5)^2/16=1 A. (-3,-4) B. (1,-5) C. (3,4) D. (5,-1)

OpenStudy (anonymous):

this you do with your eyes, not computation necessary \[\frac{(x-h)^2}{a^2}+\frac{(y-k)^2}{b^2}=1\] it the equation for an ellipse with center \((h,k)\)